OpenStudy (anonymous):

A particle starts from rest and moves in a straight line with constant acceleration. in a certain 4 seconds of its motion it travels 12m and in the next 15 seconds it travels 30m. whats the acceleration??? whats the velocity of the particle at the start of the timing then?? and whats distance it would have travelled before the timing started??? thanks

OpenStudy (anonymous):

Did you write this right? If the acceleration is constant, then the average velocity is the velocity at t = 1/2 the time interval\[v _{avg 1}=d/t=12m/4s=3m/s\]This is the instantaneous speed two seconds into the interval\[v _{avg2}=d/t=30m/15s=2m/s\]This is the instantaneous speed 7.5 seconds into the second time interval\[a=\frac {v _{avg2}-v _{avg1}} {(2s+7.5s)}=\frac {2m/s-3m/s} {(9.5s)}=-0.105m/s^{2}\]A negative acceleration means the object is slowing down.... but the problem says it starts from rest and speeds up... Something's wrong :-(